OpenStudy (anonymous):

Identify the vertex for the graph of y = 3x2 + 12x + 5. A.)(2, 41) B.)(2, 35) C.)(−2, −31) D.) (−2, −7)

OpenStudy (anonymous):

First you can change this into vertex form.

OpenStudy (saifoo.khan):

Or you can use the shortcut formula..

OpenStudy (anonymous):

True that. :D

OpenStudy (anonymous):

\[y = ax ^{2}+bx+c\] value for x coordinate vertex is given by: x = -b/2a x = -12/2*3 x = -12/6 x = -2 for y-value substitute in equation: y = 3*(-2)^2-24+5 y = -7 vertex (-2,-7)

OpenStudy (anonymous):

a(x-(-b/2a))^2+c-(b^2/4a)

OpenStudy (anonymous):

So yea the vertex turns out to be what you're doing :L

OpenStudy (anonymous):

thank you all! ♥